OpenStudy (anonymous):

write and equality and solve a number decreased by 8 is less than or equal to 14 what is the solution set

OpenStudy (anonymous):

n is the number

Parth (parthkohli):

'A number decreased by 8' will translate to \(n - 8\). 'Less than or equal to 14' will translate to \(\le 14\)

Parth (parthkohli):

The inequality will be \(n - 8 \le 14\). Can you solve it?

OpenStudy (anonymous):

is it n >22

Parth (parthkohli):

Yeah, but it's \(\le\) instead of \(>\).